# Rethinking Individual Fairness in Deepfake Detection

### v1

@@ 6.2. Results
Performance on Intra-domain Testing. We first evaluate the performance of our method under the intra-domain setting, where each model is trained and tested on the same dataset. As shown in Table 1, our method consistently outperforms both Ori and Naive across all datasets in terms of individual fairness. For example, on the FF++ dataset, our method improves the AUC by 5.77% over Ori and reduces the individual fairness loss $\mathcal{L}_{ind}^{*}$ by 0.7763. Compared to Naive, it achieves a 5.41% improvement in AUC and a 0.6373 reduction in $\mathcal{L}_{ind}^{*}$. These consistent reductions in individual fairness across all five datasets—including AI-Face (Lin et al., 2024d), which contains diffusion model-generated faces—demonstrate the effectiveness of our method in enhancing individual fairness.
@@
Plug-and-Play Flexibility Across Different Detectors. To validate the broad applicability of our framework, we integrate it into a wide range of state-of-the-art deepfake detectors across spatial-based, frequency-based, and fairness-enhanced categories. Since anchor learning may not be compatible with some detectors due to architectural constraints (e.g., UCF (Yan et al., 2023a) and PG-FDD (Lin et al., 2024c)), we apply only the semantic-agnostic individually fair learning in these methods. As demonstrated in Table 3, regardless of the different detection strategies, our method successfully and consistently improves individual fairness and detection utility. For example, when applied to SRM (Luo et al., 2021), a frequency-based model, our method reduces $\mathcal{L}_{ind}^{*}$ from 0.5797 to 0.3688 and improves AUC from 96.29% to 97.06%. Particularly notable is the improvement on PG-FDD (Lin et al., 2024c), where $\mathcal{L}_{ind}^{*}$ is lowered by 0.1387, demonstrating the compatibility of our framework with existing fairness-focused solutions. These results confirm that our method is model-agnostic and can be easily integrated to enhance individual fairness across diverse deepfake detection strategies.
@@
PG-FDD (Lin et al., 2024c) 97.59 0.8538 0.6063
Fairness- enhanced Ours (PG-FDD) 97.93 0.7022 0.4676
@@ 7. Conclusion
In this paper, we challenge the conventional understanding of individual fairness in the context of deepfake detection. We are the first to identify a fundamental contradiction: the original principle of individual fairness fails in deepfake detection due to the high semantic similarity between real and fake samples. To address this issue, we propose a novel and generalizable framework that integrates seamlessly with existing detectors to improve individual fairness without sacrificing detection utility. Our approach introduces two key components: (1) anchor learning to enhance fairness generalization, and (2) semantic-agnostic individual fair learning to break high-level semantic correlations and expose manipulation artifacts. Extensive experiments across various datasets and detector backbones demonstrate the effectiveness of our proposed method.
@@
Future Work. Although the AI-Face dataset is used in our experiments because it contains deepfake video-based images exhibiting semantic similarity issues, the effectiveness of our method on detecting faces generated solely by GANs or diffusion models remains unclear. These generated faces typically differ significantly from real ones and may not exhibit the semantic similarity challenges discussed in this work. How do we improve individual fairness in such a scenario? We plan to investigate this problem in future research. We also plan to extend our methods to improve individual fairness in detecting AI-generated voices and even multi-modal content such as deepfake videos.
@@ 8. Algorithm
15:        Compute gradient approximation for Eq. (6)
@@
20:Output: $\theta_{l^{*}}$, where $l^{*}$ is the best iterate satisfying Eq. (6) with the lowest objective.
@@ 9.1. Description of The Datasets
Here, we show the total number of training and test samples for each dataset included in our experiments, as in Table 6. While our method does not inherently require the demographic attributes (i.e., gender, race) included in the datasets, we leveraged these annotations to benchmark our proposed method against fairness-aware baselines such as DAW-FDD (Ju et al., 2024) and PG-FDD (Lin et al., 2024c), which explicitly incorporate these features into their pipelines.
@@ 9.3. Implementation Details
FF++ (Rossler et al., 2019) DFDC (Brian et al., 2020) Celebdf-DF (Li et al., 2020) DFD (Google and Jigsaw, 2019) AI-Face (Lin et al., 2024d)

### v2

@@ 6.2. Results
Performance on Intra-domain Testing. We first evaluate the performance of our method under the intra-domain setting, where each model is trained and tested on the same dataset. As shown in Table 1, our method consistently outperforms both Ori and Naive across all datasets in terms of individual fairness. For example, on the FF++ dataset, our method improves the AUC by 5.77% over Ori and reduces the individual fairness loss $\mathcal{L}_{ind}^{*}$ by 0.7763. Compared to Naive, it achieves a 5.41% improvement in AUC and a 0.6373 reduction in $\mathcal{L}_{ind}^{*}$. These consistent reductions in individual fairness across all five datasets—including AI-Face (Lin et al., 2024e), which contains diffusion model-generated faces—demonstrate the effectiveness of our method in enhancing individual fairness.
@@
Plug-and-Play Flexibility Across Different Detectors. To validate the broad applicability of our framework, we integrate it into a wide range of state-of-the-art deepfake detectors across spatial-based, frequency-based, and fairness-enhanced categories. Since anchor learning may not be compatible with some detectors due to architectural constraints (e.g., UCF (Yan et al., 2023a) and PG-FDD (Lin et al., 2024d)), we apply only the semantic-agnostic individually fair learning in these methods. As demonstrated in Table 3, regardless of the different detection strategies, our method successfully and consistently improves individual fairness and detection utility. For example, when applied to SRM (Luo et al., 2021), a frequency-based model, our method reduces $\mathcal{L}_{ind}^{*}$ from 0.5797 to 0.3688 and improves AUC from 96.29% to 97.06%. Particularly notable is the improvement on PG-FDD (Lin et al., 2024d), where $\mathcal{L}_{ind}^{*}$ is lowered by 0.1387, demonstrating the compatibility of our framework with existing fairness-focused solutions. These results confirm that our method is model-agnostic and can be easily integrated to enhance individual fairness across diverse deepfake detection strategies.
@@
PG-FDD (Lin et al., 2024d) 97.59 0.8538 0.6063
Fairness-
enhanced
Ours (PG-FDD) 97.93 0.7022 0.4676
@@ 7. Conclusion
In this paper, we challenge the conventional understanding of individual fairness in the context of deepfake detection. We are the first to identify a fundamental contradiction: the original principle of individual fairness fails in deepfake detection due to the high semantic similarity between real and fake samples. To address this issue, we propose a novel and generalizable framework that integrates seamlessly with existing detectors to improve individual fairness without sacrificing detection utility.
@@
Future Work. Although the AI-Face dataset is used in our experiments because it contains deepfake video-based images exhibiting semantic similarity issues, the effectiveness of our method on detecting faces generated solely by GANs or diffusion models remains unclear. These generated faces typically differ significantly from real ones and may not exhibit the semantic similarity challenges discussed in this work. How do we improve individual fairness in such a scenario? We plan to investigate this problem in future research. We also plan to extend our methods to improve individual fairness in detecting AI-generated voices.
@@ Acknowledgements.
8. Related Work
Deepfake Detection. Current deepfake detection methods can be categorized into three primary groups based on the features they employ. The first category hinges on identifying inconsistencies in the physical and physiological characteristics of deepfakes (Lin et al., 2024c). For example, inconsistent corneal specular highlights (Hu et al., 2021), the irregularity of pupil shapes (Guo et al., 2022a, b), eye blinking patterns (Li et al., 2018), eye color difference (Matern et al., 2019), facial landmark locations (Yang et al., 2019), etc. The second category concentrates on signal-level artifacts introduced during the synthesis process, especially those from the frequency domain (Qian et al., 2020). These methods encompass various techniques, such as examining disparities in the frequency spectrum (Khayatkhoei and Elgammal, 2022; Dzanic et al., 2020) and utilizing checkerboard artifacts introduced by the transposed convolutional operator (Frank et al., 2020; Zhang et al., 2019). However, the methods in the above two categories usually exhibit relatively low detection performance. Therefore, the largest portion of existing detection methods fall into the data-driven category, including (Guo et al., 2022c; Pu et al., 2022a; Yan et al., 2023a; Liu et al., 2024; Hong et al., 2024b; Choi et al., 2024). These methods leverage various types of DNNs trained on both authentic and deepfake videos to capture specific discernible artifacts. While these methods achieved strong detection utility, they overlooked detection fairness, resulting in large performance disparities among groups (e.g., gender).
Fairness in Deepfake Detection. Many studies (Trinh and Liu, 2021; Hazirbas et al., 2021; Pu et al., 2022b; Xu et al., 2024; Lin et al., 2024e; Nadimpalli and Rattani, 2022) have identified fairness issues in deepfake detection. For instance, Trinh et al. (Trinh and Liu, 2021) found biases present in both datasets and detection models, resulting in significant disparities in error rates across demographic groups. Similar findings were reported by Hazirbas et al. (Hazirbas et al., 2021). Pu et al. (Pu et al., 2022b) evaluated the fairness of the MesoInception-4 model using the FF++ dataset, observing performance bias related to gender. Nadimpalli et al. (Nadimpalli and Rattani, 2022) also highlighted substantial biases and proposed a preprocessing method involving a gender-balanced dataset to mitigate gender-based performance disparities. However, this approach delivered only modest improvements and necessitated extensive data annotation efforts. To further support fairness-oriented algorithm development, (Xu et al., 2024) and (Lin et al., 2024e) performed comprehensive analyses of bias in deepfake detection, augmenting datasets with diverse demographic annotations. Recent works by Ju et al. (Ju et al., 2024) and Lin et al. (Lin et al., 2024d) have subsequently concentrated on enhancing group-level fairness through algorithmic interventions. However, these studies have not comprehensively examined fairness at the individual level with solutions, an aspect that remains largely unaddressed in recent deepfake detection research and constitutes the primary focus of our work.
@@ 9. Algorithm
15:   Compute gradient approximation for $\min_{\theta}\mathcal{L}\textbf{(}\theta+\epsilon^{*}\textbf{)}\end{aligned}$
@@
20:Output: $\theta_{l^{*}}$, where $l^{*}$ is the best iterate satisfying $\min_{\theta}\mathcal{L}\textbf{(}\theta+\epsilon^{*}\textbf{)}\end{aligned}$ with the lowest objective.
@@ 10.1. Description of The Datasets
Here, we show the total number of training and test samples for each dataset included in our experiments, as in Table 6. While our method does not inherently require the demographic attributes (i.e., gender, race) included in the datasets, we leveraged these annotations to benchmark our proposed method against fairness-aware baselines such as DAW-FDD (Ju et al., 2024) and PG-FDD (Lin et al., 2024d), which explicitly incorporate these features into their pipelines.
@@ 10.3. Implementation Details
FF++ (Rossler et al., 2019) DFDC (Brian et al., 2020) Celebdf-DF (Li et al., 2020) DFD (Google and Jigsaw, 2019) AI-Face (Lin et al., 2024e)